We need to spend $20,000 per year for 3 years on new restaurant equipment. How much...

  1. We need to spend $20,000 per year for 3 years on new restaurant equipment. How much do we need to have in an account to afford this if we earn 5% on our investments?
    •   __________________
    •   20,000  _________
    •   _______   N
    •   _______  1/Y
    •   CPT _______
    •   Answer: _____________
Answer #1
Amount to be in account = Annual cash flow * Present value of annuity of 1
= $ 20,000.00 * 2.723248
= $ 54,464.96
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.05)^-3)/0.05 i = 5%
= 2.72324803 n = 3